Steckly Company had the following balances on its balance sheet at Beginning of Year2. Both are positive equity account balances.

Retained earnings $42,000

Accumulated other comprehensive income $4,000

For Year 2 Steckly reported net income of $23,000. They also reported Comprehensive income of $26,500. No dividends were declared or paid.

1: What was "Other Comprehensive income" during Year2?

2: What will be the balance in Retained earnings on the Year2 balance sheet?

3: What will be the balance in Accumulated other comprehensive income on the Year2 balance sheet?
Answer #1

1) Other Comprehensive income = Comprehensive income-Net income = 26500-23000 = 3500

2) Ending retained earnings Year 2 = 42000+23000 = 65000

3) Ending accumulated other comprehensive income = 4000+3500 = 7500